# Are Existing Road Design Guidelines Suitable for Autonomous Vehicles?

### v1

@@ 3.1 Generating Natural Scenarios (C1)
$rule2$ The fixed position attack objects (OG Boxes) should be at least 0.6 meters away from the main road. The main road refers to a road with at least two lines.
@@
$rule3$ The fixed position attack objects (OG Boxes) cannot be put near the minor road. A minor road refers to a road with one line.
@@
$rule4$ The splay corner of the entrance culvert, bin centre access, substation access, MDF room access, and fire engine access should be at least 1m to Palm, 1.5m from a small to medium tree, and 2.5m from a large tree.
@@
$rule5$ The scupper pipe/drain should be at least 1m from a Palm, 1.5m from a small to medium tree and 2.5m from a large tree.
$tree.h<1.5\implies dis(tree,pipe)>1\land(tree.h\geq 1.5\land tree.h<5)\implies dis(tree,pipe)>1.5\land tree.h\geq 5\implies dis(tree,pipe)>2.5$
$rule6$ The trees should be at least 3m from a Lamp post.
@@
$rule7$ The fixed position attack objects (OG Boxes), TAS manhole, sewer line, manhole, electrical post, fire hydrant, SCV box, lighting control box, traffic control box, traffic lights, or lightning pits should be at least 2m to Palm or small to medium trees, and 2.5m to large trees.
$tree.h<5\implies dis(tree,obj.fix)>2\land tree.h\geq 5\implies dis(tree,obj.fix)>2.5$
$rule8$ Trees should be at least 3m from the edge of a footpath crossing.
@@
$rule9$ Trash bins should not block the road or footpath.
@@
$rule10$ Trash bins should maintain a minimum distance of 0.5 meters from each other and other objects like mailboxes, walls, lamp posts, utility boxes, vehicles, etc.
@@
$rule11$ Wheels and handles of trash bins should be placed away from the road.
@@
$rule12$ Make sure nothing is in between the trash bins and the road.
@@
As mentioned, we overcome C1 by utilising existing regulations and guidelines for the placement of various roadside objects. In particular, we refer to government documents including the Code of Practice for Works on Public Streets (COP) [5] and Guidelines on Greenery Provision and Tree Conservation (GGPTC) [7]. The COP, formulated by Singapore’s Land Transport Authority, delineates the necessary procedures and regulations for conducting activities on public roads, whereas the GGPTC, introduced by Singapore’s National Parks Board, governs the placement and preservation of tress within urban landscapes. For the placement of trash bins, we refer to the guidelines provided by three prominent recycling companies: HDS, KIMBLE, and GRANGER [8, 9, 10].
We formalise the notion of natural scenario by translating the regulations/guidelines described in these documents into mathematical formulas. Though in natural language, the guidelines are quite specific, allowing for a relatively direct translation of the rules to formulas. Table I lists these 12 specifications along with some descriptions. Specifications 1–3 refer to section 6.5 of the COP [5]. Specifications 4–8 refer to pages 95-97 of the GGPTC [7]. Specifications 9–11 pertain to key aspects of the guidelines provided by the waste recycling companies [8, 9, 10]. Note that while we have focused on guidelines from one country to ensure consistency between our rules, the general approach of formalising rules can be applied to the guidelines of any other country or jurisdiction.
In our rules, the function $dis(a,b)$ returns the distance between two objects, $a$ and $b$, while the function $lanenum(road)$ returns the count of lanes within the designated $road$. Each distinct category of objects is denoted by a class variable. For example, the class $tree$ serves as a general representation for trees and includes a characteristic attribute, $tree.h$, corresponding to height.
It is crucial to have a diverse range of attack objects at our disposal. Several simulators, including AWSIM [35] and LGSVL [19]), are built upon the Unity Real-Time Development Platform [36]. The Unity Asset Store [20] has a vast library of 3D asset packages, with an open-door policy that allows anyone to contribute their creations. Each asset package provides necessary elements, such as meshes of 3D objects and 2D textures, for constructing objects inside the simulators. We can access thousands of objects that are commonly encountered during driving such as rocks, trash bins, and hydrants for free. Hence, our strategy involves exploring Unity’s asset library [20] to discover suitable adversarial objects.
@@
3.2 Greedy Gradient-Guided Black Box Search (C2)
We require an efficient algorithm for identifying suitable attack scenarios in a black box setting. Given our objective is to induce ADS behaviours that violate traffic law specifications, it is important that the algorithm can uncover different ways to violate the specification. For instance, consider the specification $speed>10\land speed<100$. There are two distinct ways to violate it: first by inducing $speed$ to a value smaller than 10, and second by inducing it to a value larger than 100.
In this section, we present the detailed design of our algorithm. Given the time-consuming nature of simulator-based evaluations for ADSs, our primary goal is to attain convergence as fast as possible. To achieve this, our algorithm is ‘greedy’, in the sense of initially exploring adversarial scenarios that appear most likely to violate the given specification, and is also ‘gradient guided’, in the sense that it systematically expands its exploration based on these initial findings.
We present the algorithm in three parts. First, we delve into the encoding of (natural) scenarios as vectors, which forms the foundation of subsequent steps (e.g. mutation). Second, we present the design of our objective function which provides a quantified way to sufficiently ‘cover’ different ways of violating a specification, ensuring diversity in the results. Finally, we introduce our new greedy gradient-guided search algorithm, which forms the core of the approach.
3.2.1 Encoding of Scenarios
We represent each adversarial scenario as a vector $\phi$ with dimensions $n\times d$. Here, $n$ refers to the number of adversarial objects, while $d$ corresponds to the four essential dimensions of each adversarial object: forward, right, rotation, and type. The forward and right dimensions indicate the relative position of the adversarial object with respect to the start position of the ego vehicle. For example, if the value for forward is $10$, and the value for right is $-2$, the object is located at the left front of the ego vehicle. The rotation dimension represents the placement angle of the adversarial object. The type is a sequence number corresponding to the adversarial objects library, allowing us to easily identify and categorise different types of adversarial objects. In addition, each scenario also provides the initial and destination positions for the ADS journey.
@@ 3.2.1 Encoding of Scenarios
3.2.2 Coverage-based Objective Function Design
@@
Let $\Theta$ be a set of property specifications and $\Lambda$ a set of adversarial scenarios. Our optimisation problem is:
@@
We thus require a set of adversarial scenarios that minimises the robustness function across several property specifications. Intuitively, the robustness function reflects the distance to a specification violation, meaning it measures how close the current trace $\pi$ is to violating the specification. Firstly, we must establish a precise definition for the fundamental objective function, denoted as $\rho(\phi,\pi)$. Secondly, we require a systematic approach to replace the specification with a set of smaller formulas $(\Theta)$, each one characterising a different way to violate the original specification. Note that we adapt the quantitative semantics and decomposition from LawBreaker [17].
Quantitative Semantics. We define our objective function using the quantitative semantics of STL [33, 38, 39], which produces a numerical robustness degree.
@@
$\rho(\mu,\pi,t)=\begin{cases}-\pi_{t}(f(x_{0},x_{1},\cdots,x_{k}))&\text{if $\sim$ is $\leq$ or $<$}\\ \pi_{t}(f(x_{0},x_{1},\cdots,x_{k}))&\text{if $\sim$ is $\geq$ or $>$}\\ \mid\pi_{t}(f(x_{0},x_{1},\cdots,x_{k}))\mid&\text{if $\sim$ is $\neq$}\\ -\mid\pi_{t}(f(x_{0},x_{1},\cdots,x_{k}))\mid&\text{if $\sim$ is $=$}\end{cases}$
where $t$ is the time step and $\pi_{t}(e)$ is the valuation of expression $e$ at time $t$ in $\pi$.
@@
Example 3.1.
Let $\varphi=\Box(speed<100)$, i.e. the speed limit is $100$km/h. Suppose $\pi$ is $\langle(speed\mapsto 0,\dots),(speed\mapsto 0.5,\dots),\cdots(speed\mapsto 90,\dots)\rangle$, where the ego vehicle’s max $speed$ is $90$km/h at the last time step. We have $\rho(\varphi,\pi)=\rho(\varphi,\pi,0)=min_{t\in[0,|\pi|]}(100-\pi_{t}(speed))=10$. This means that trace $\pi$ satisfies $\varphi$, and the robustness value is 10.
Deriving ‘Violation Goals’ from a Specification. We transform a specification into a set of smaller formulas, such that the violation of any one formula implies the violation of the original specification. The rationale is that this will allow us to evaluate the robustness degree with respect to multiple different ways of violating the original specification, rather than just focusing on the ‘easiest’ one. Here, we adopt the methodology of LawBreaker [17], which defines a method to derive a set of violation goals based on the given complex specification $\phi$. We write $\Theta(\phi)$ to denote a set of ‘violation goals’ of $\phi$, i.e. that satisfy the following proposition:
Proposition 3.2.
@@
$\Theta(\phi_{1}\land\phi_{2})=\Theta(\phi_{1})\cup\Theta(\phi_{1});\Theta(\bigcirc\phi)=\{\bigcirc x~{}|~{}x\in\Theta(\phi)\};$
@@
$n(\mu)=\begin{cases}e_{1}\geq(resp.>)e_{2}&\text{if $\mu$ is $e_{1}<(resp.\leq)e_{2}$};\\ e_{1}\leq(resp.<)e_{2}&\text{if $\mu$ is $e_{1}>(resp.\geq)e_{2}$}\\ e_{1}\neq(resp.=)e_{2}&\text{if $\mu$ is $e_{1}=(resp.\neq)e_{2}$}\\ \end{cases}$
Propositions 3.2 and 4 can be proved by structural induction. We refer the readers to [17, 40] for the detailed proofs.
Example 3.4.
The value of $\Theta(\phi)$ is calculated recursively. For example, given a specification $\phi=\Box(\mu_{1}\land\mu_{2})$, we can get $\Theta(\phi)$ as follows:
@@
    First step is to deal with calculations of primitive elements: $\Theta(\mu_{1})=\{\mu_{1}\},\Theta(\mu_{2})=\{\mu_{2}\}$
@@
i.e. there exist two violation goals for $\phi$, and the violation of either one of them can result in the violation of $\phi$.
3.2.3 Greedy Gradient Guided Query
With the derivation of violation goals reviewed, we now propose our new greedy gradient-guided query algorithm. We first describe the gradient calculation process before presenting the algorithm as a whole.
Gradient Calculation. In our black-box setting, gradient calculations are important for guiding the generation of adversarial scenarios. These gradients are constructed by differentiating between the original scenario and the scenario following the modifications. We define the gradient calculation below:
@@ 3.2.3 Greedy Gradient Guided Query
Let $\phi$ be a property specification, $\lambda$ be the initial scenario, $\lambda^{\prime}$ be the scenario after modification, and $E(\lambda)$/$E(\lambda^{\prime})$ denote the encoding of scenario $\lambda$/$\lambda^{\prime}$. The gradient calculation is formally defined as follows:
@@
where $\pi$ and $\pi^{\prime}$ correspond to the traces of the ADS in response to scenarios $\lambda$ and $\lambda^{\prime}$, respectively.
@@
Input: $\phi$, $n$ (number of adversarial objects), and $M$ (maximal queries)
@@ 1 Let $\Theta_{r}=\Theta(\phi)$ be the set of uncovered formulas in $\Theta(\phi)$;
2 Let $seed$ be a pair of a gradient value and a mutable element; initially the gradient value $seed.gradient$ is 0 and the mutable element $seed.element$ is $Null$;
@@ 12                         Remove $\varphi$ from $\Theta_{r}$; Add $s_{k}$ into $\Gamma$;
16                         $seed.gradient=G(\pi_{0},\pi_{k},\varphi)$;
@@
24      Generate the next $s_{0}$ based on $seed$;
@@
Algorithm 1 The $\mathtt{TrashFuzz}$ Algorithm
Overall Algorithm. The overall algorithm is shown in Algorithm 1. First, we prepare a random scenario as our initial test case and calculate the subdivisions of the user-given specification $\varphi$ to obtain $\Theta_{r}$, then initialize $seed$ with $seed.gradient$ value 00 and $seed.element$ value $Null$. Note that $seed$ is a pair storing gradient value and a corresponding scenario.
Next, using the initial adversarial scenario $\lambda_{0}$, we execute it to generate the trace $\pi_{0}$. The adversarial scenario contains several adversarial objects, and for each of these objects, we apply random mutations to create new scenarios. Note that the mutations have been designed to ensure that the ADS remains compliant with the rules listed in Table I, as well as with other physical constraints such as the requirement for objects to be grounded. For each newly generated scenario, we also execute it to produce a trace and assess it against all the remaining specifications in set $\Theta_{r}$. We remove $\varphi$ from the set if it is violated, i.e. if $\rho(\xi,\pi_{i})\leq 0$, and add the corresponding scenario $\lambda_{i}$ to the output set $\Gamma$. After executing and processing all the cases generated by the initial test case, we utilise the gradient calculation defined above to choose the seed for the next round. In essence, we designate the case with the largest absolute gradient value as the seed for the subsequent round.
@@
4 Implementation and Evaluation
In this section, we present our implementation of $\mathtt{TrashFuzz}$ for the Apollo ADS and evaluate its effectiveness in identifying natural scenarios that lead to traffic law violations, as well as other key aspects. For additional experimental materials and results, we direct readers to our website [37].
4.1 Implementation
Implementing $\mathtt{TrashFuzz}$ involves three broad steps. First, some basic integration with and between the existing ADS and simulator, including a communication bridge, a trace generator, and the means of both checking a specification as well as calculating its robustness with respect to a trace. Second, the generation of natural scenarios consisting of (benign-appearing) adversarial roadside objects, as detailed in Section 3.1. Finally, the implementation of our greedy gradient-based fuzzing engine, as elaborated on in Section 3.2.
The target of our evaluation is version 7.0 of the Apollo ADS [18] (the most recent version at the time of experimentation) with the LGSVL [19] simulator. To establish communication between Apollo and LGSVL, we utilise the official communication bridge, CyberRT [41]. For trace generation and specification evaluation/robustness, we use the implementation strategy described by LawBreaker [17]. In particular, the trace generator creates a trace $\pi$ by processing messages subscribed to by the ADS, whereas for specifying traffic laws, we utilise the domain-specific language provided by LawBreaker. To implement the quantitative semantics of those traffic laws, we embedded the RTAMT tool [39] to compute the robustness of the specifications with respect to the trace obtained from the bridge. (Note that implementing $\mathtt{TrashFuzz}$ for Autoware/CARLA follows a largely similar strategy, except for substituting the ROS bridge [42] for CyberRT.)
In order to generate natural scenarios, we sourced common roadside objects from the Unity Asset Store [20]. As LGSVL is a Unity-based simulator, these could be directly placed into scenarios. The roadside objects we utilised in $\mathtt{TrashFuzz}$ have been assembled into a single library that we have made available online [37]. To ensure the objects are placed according to conventional road design guidelines, we constrained their placement according to the functions in Table I. These functions take a scenario as input and return a Boolean value to indicate whether the generated scenario is valid or not according to the guidelines.
Finally, we integrated the above components and implemented our greedy gradient-based fuzzing engine as described in Section 3.2. Our experiments were conducted using two machines, each equipped with 32GB of memory, an Intel i7-10700k CPU, and an RTX 2080Ti graphics card. These two machines run on Linux (Ubuntu 20.04.5 LTS) and Windows (Windows 10 Pro) operating systems, respectively.
4.2 Evaluation
We conducted experiments to answer four Research Questions (RQs).
  • RQ1: Can $\mathtt{TrashFuzz}$ produce valid scenarios?

### v3

@@ III-A Generating Natural Scenarios (C1)
$rule2$ The fixed position attack objects (OG boxes) should be at least 0.6 m away from the main road (2 lanes).
@@
$rule3$ The fixed position attack objects cannot be near a minor road (1 lane).
@@
$rule4$ The splay corner of the entrance culvert, bin centre access, substation access, MDF room access, and fire engine access must be at least 1 m from palms, 1.5 m from small/medium trees, and 2.5 m from large trees.
@@
$rule5$ A scupper pipe/drain must be at least 1 m from palms, 1.5 m from small/medium trees, and 2.5 m from large trees.
$(tree.h<1.5\implies dis(tree,pipe)>1)\land(tree.h\geq 1.5\land tree.h<5\implies dis(tree,pipe)>1.5)\land(tree.h\geq 5\implies dis(tree,pipe)>2.5)$
$rule6$ Trees must be at least 3 m from a lamp post.
@@
$rule7$ Fixed objects (e.g., OG box, manhole, SCV box, lighting box, traffic light) must be at least 2 m from small/medium trees, and 2.5 m from large trees.
$(tree.h<5\implies dis(tree,obj.fix)>2)\land(tree.h\geq 5\implies dis(tree,obj.fix)>2.5)$
$rule8$ Trees must be at least 3 m from the edge of a footpath crossing.
@@
$rule9$ Trash bins should not block roads or footpaths.
@@
$rule10$ Trash bins must be at least 0.5 m from each other and from mailboxes, walls, lamp posts, utility boxes, vehicles, etc.
@@
$rule11$ Wheels/handles of bins must be oriented away from the road.
@@
$rule12$ Ensure no object lies between trash bins and the road.
@@
As mentioned, we overcome C1 by utilising existing regulations and guidelines for the placement of various roadside objects. In particular, we refer to government documents including the Code of Practice for Works on Public Streets (COP) [29] and Guidelines on Greenery Provision and Tree Conservation (GGPTC) [36]. The COP, formulated by Singapore’s Land Transport Authority, delineates the necessary procedures and regulations for conducting activities on public roads, whereas the GGPTC, introduced by Singapore’s National Parks Board, governs the placement and preservation of trees within urban landscapes. For the placement of trash bins, we refer to the guidelines provided by three prominent recycling companies: HDS, KIMBLE, and GRANGER [22, 26, 16].
We formalise the notion of natural scenario by translating the regulations/guidelines described in these documents into mathematical formulas. Though in natural language, the guidelines are quite specific, allowing for a relatively direct translation of the rules to formulas. Table I lists these 12 specifications that jointly define natural scenarios, all of which must be satisfied. Specifications 1–3 refer to section 6.5 of the COP [29]. Specifications 4–8 refer to pages 95-97 of the GGPTC [36]. Specifications 9–11 pertain to key aspects of the guidelines provided by the waste recycling companies [22, 26, 16]. Note that while we have focused on guidelines from one country to ensure consistency between our rules, the general approach of formalising rules can be applied to the guidelines of any other country or jurisdiction.
In our rules, the function $dis(a,b)$ returns the distance between two objects, $a$ and $b$, while the function $lanenum(road)$ returns the count of lanes within the designated roadroad. Each distinct category of objects is denoted by a class variable. For example, the class treetree serves as a general representation for trees and includes a characteristic attribute, $tree.h$, corresponding to height.
It is crucial to have a diverse range of attack objects at our disposal. Several simulators, including AWSIM [1] and LGSVL [44], are built upon the Unity Real-Time Development Platform [55]. The Unity Asset Store [56] has a vast library of 3D asset packages, with an open-door policy that allows anyone to contribute their creations. Each asset package provides necessary elements, such as meshes of 3D objects and 2D textures, for constructing objects inside the simulators. We can access thousands of objects that are commonly encountered during driving such as rocks, trash bins, and hydrants for free. Hence, our strategy involves exploring Unity’s asset library [56] to discover suitable adversarial objects.
@@
III-B Greedy Gradient-Guided Black-Box Search (C2)
We require an efficient algorithm for identifying attack scenarios in a black-box setting. Given our objective is to induce ADS behaviours that violate traffic law specifications, it is important that the algorithm can uncover different ways to violate the specification. For instance, consider the specification $(speed>10)\wedge(speed<100)$. There are two distinct ways to violate it: first by inducing speedspeed to a value smaller than 10, and second by inducing it to a value larger than 100.
In this section, we present the detailed design of our algorithm. Given the time-consuming nature of simulator-based evaluations for ADSs, our primary goal is to attain convergence as fast as possible. To achieve this, our algorithm is greedy, in the sense of initially exploring adversarial scenarios that appear most likely to violate the given specification, and is also gradient guided, in the sense that it systematically expands its exploration based on these initial findings.
We present the algorithm in three parts. First, we delve into the encoding of (natural) scenarios as vectors, which forms the foundation of subsequent steps (e.g., mutation). Second, we present the design of our objective function which provides a quantified way to sufficiently cover different ways of violating a specification, ensuring diversity in the results. Finally, we introduce our new greedy gradient-guided search algorithm, which forms the core of our approach.
III-B1 Encoding of Scenarios
We represent each adversarial scenario as a vector $\phi$ with dimensions $n\times d$. Here, $n$ refers to the number of adversarial objects, while $d$ corresponds to the four essential dimensions of each adversarial object: forward, right, rotation, and type. The forward and right dimensions indicate the relative position of the adversarial object with respect to the start position of the ego vehicle. For example, if the value for forward is $10$, and the value for right is 2-2, the object is located at the left front of the ego vehicle. The rotation dimension represents the placement angle of the adversarial object. The type is a sequence number corresponding to the adversarial objects library, allowing us to easily identify and categorise different types of adversarial objects. In addition, each scenario also provides the initial and destination positions for the ADS journey.
@@ III-B1 Encoding of Scenarios
III-B2 Coverage-based Objective Function Design
@@
Let $\Theta$ be a set of traffic law specifications and $\Lambda$ a set of adversarial scenarios. Our optimisation problem is:
@@
We thus require a set of adversarial scenarios that minimises the robustness function across several property specifications. Intuitively, the robustness function reflects the distance to a specification violation, meaning it measures how close the current trace $\pi$ is to violating the specification. First, we must establish a precise definition for the fundamental objective function, denoted as $\rho(\phi,\pi)$. Second, we require a systematic approach to replace the specification with a set of smaller formulas $(\Theta)$, each one characterising a different way to violate the original specification.
We define our objective function using the quantitative semantics of STL [33, 13, 37], which produces a numerical robustness degree.
@@
[…]
where $t$ is the time step and $\pi_{t}(e)$ is the valuation of expression $e$ at time $t$ in $\pi$;
@@
Example III.1.
Let $\varphi=\Box(speed<100)$, i.e., the vehicle’s speed must always remain below $100$ km/h. Suppose $\pi$ is $\langle(speed\mapsto 0,\dots),(speed\mapsto 0.5,\dots),\cdots(speed\mapsto 90,\dots)\rangle$, where the ego vehicle’s max speedspeed is 9090km/h at the last time step. We have $\rho(\varphi,\pi)=\rho(\varphi,\pi,0)=min_{t\in[0,|\pi|]}(100-\pi_{t}(speed))=10$. This means that trace $\pi$ satisfies $\varphi$, and the robustness value is 10.
Next, we transform a specification into a set of smaller formulas, such that the violation of any one formula implies the violation of the original specification. The rationale is that this will allow us to evaluate the robustness degree with respect to multiple different ways of violating the original specification, rather than just focusing on the ‘easiest’ one. Here, we adopt the methodology of LawBreaker [50], which defines a method to derive a set of violation goals based on the given complex specification $\phi$. We write $\Theta(\phi)$ to denote a set of ‘violation goals’ of $\phi$, i.e., that satisfy the following proposition:
Proposition III.2.
@@
$\Theta(\phi_{1}\land\phi_{2})=\Theta(\phi_{1})\cup\Theta(\phi_{2});\Theta(\bigcirc\phi)=\{\bigcirc x~|~x\in\Theta(\phi)\};$
@@
[…]
Propositions III.2 and III.3 can be proved by structural induction. We refer the readers to [50] for the detailed proofs.
Example III.4.
The value of $\Theta(\phi)$ is calculated recursively. For example, given a specification $\phi=\Box(\mu_{1}\land\mu_{2})$, we can obtain $\Theta(\phi)$ as follows:
@@
    First, we apply $\Theta$ to the primitive elements: $\Theta(\mu_{1})=\{\mu_{1}\},\Theta(\mu_{2})=\{\mu_{2}\}$
@@
i.e., there exist two violation goals for $\phi$, and the violation of either one of them can result in the violation of $\phi$.
III-B3 Greedy Gradient Guided Query
With the derivation of violation goals reviewed, we now propose our novel greedy gradient-guided query algorithm. We first describe the gradient calculation process before presenting the algorithm as a whole.
In our black-box setting, gradient calculations are important for guiding the generation of adversarial scenarios. These gradients are constructed by differentiating between the original scenario and the scenario resulting from modifications. We define the gradient calculation below:
@@ III-B3 Greedy Gradient Guided Query
Let $\phi$ be a property specification, $\lambda$ be the initial scenario, $\lambda^{\prime}$ be the scenario after modification, and $E(\lambda)$$E(\lambda^{\prime})$ denote the encoding of scenario $\lambda$$\lambda^{\prime}$. The gradient calculation is formally defined as follows:
@@
where $\pi$ and $\pi^{\prime}$ are the ADS traces under scenarios $\lambda$ and $\lambda^{\prime}$, respectively.
@@
Input: $\phi$ (traffic law), $n$ (number of adversarial objects), and $M$ (maximal queries)
@@ 1 Let $\Theta_{r}=\Theta(\phi)$ be the set of uncovered formulas in $\Theta(\phi)$;
2 Let seedseed be a pair storing a gradient value and mutable element; initially the gradient value $seed.gradient$ is 0 and the mutable element $seed.element$ is NullNull;
@@ 12           Remove $\varphi$ from $\Theta_{r}$; Add $s_{k}$ into $\Gamma$;
16           $seed.gradient=G(\lambda_{0},\lambda_{k},\varphi)$;
@@
24 Generate the next $s_{0}$ based on seedseed;
@@
Algorithm 1 The TrashFuzz Algorithm
The overall TrashFuzz algorithm is shown in Algorithm 1. First, we generate a random scenario as the initial test case and decompose the traffic law specification $\phi$ to obtain $\Theta_{r}$, then initialise seedseed with $seed.gradient$ value 0 and $seed.element$ value $Null$. Note that seedseed is a pair storing a gradient value and the corresponding mutable element.
Next, using the initial adversarial scenario $\lambda_{0}$, we execute it to generate the trace $\pi_{0}$. The adversarial scenario contains several adversarial objects, and for each of these objects, we apply random mutations to create new scenarios. Note that the mutations have been designed to ensure that the ADS remains compliant with the rules listed in Table I, as well as with other physical constraints such as the requirement for objects to be grounded. For each newly generated scenario, we also execute it to produce a trace and assess it against all the remaining specifications in set $\Theta_{r}$. We remove $\varphi$ from the set if it is violated, i.e., if $\rho(\varphi,\pi_{i})\leq 0$, and add the corresponding scenario $\lambda_{i}$ to the output set $\Gamma$. After executing and processing all the cases generated by the initial test case, we utilise the gradient calculation defined above to choose the seed for the next round. In essence, we designate the case with the largest absolute gradient value as the seed for the subsequent round.
@@
IV Implementation and Evaluation
In this section, we present our implementation of TrashFuzz for the Apollo ADS and evaluate its effectiveness in identifying natural scenarios that lead to traffic law violations, as well as other key aspects. For additional experimental materials and results, we direct readers to our website [51].
IV-A Implementation
Implementing TrashFuzz involves three broad steps. First, some basic integration with and between the existing ADS and simulator, including a communication bridge, a trace generator, and the means of both checking a specification as well as calculating its robustness with respect to a trace. Second, the generation of natural scenarios consisting of (benign-appearing) adversarial roadside objects, as detailed in Section III-A. Finally, the implementation of our greedy gradient-based fuzzing engine, as elaborated on in Section III-B.
The target of our evaluation is the LGSVL [44] simulator with version 7.0 of the Apollo ADS [3] (the latest version stable with LGSVL). To establish communication between Apollo and LGSVL, we utilise the official communication bridge, CyberRT [30]. For trace generation and specification evaluation/robustness, we use the implementation strategy described by LawBreaker [50]. In particular, the trace generator creates a trace $\pi$ by processing messages subscribed to by the ADS, whereas for specifying traffic laws, we utilise the domain-specific language provided by LawBreaker. To implement the quantitative semantics of those traffic laws, we embedded the RTAMT tool [37] to compute the robustness of the specifications with respect to the trace obtained from the bridge. (Note that implementing TrashFuzz for Autoware/CARLA would follow a largely similar strategy, except for substituting the ROS bridge [6] for CyberRT.)
In order to generate natural scenarios, we sourced common roadside objects from the Unity Asset Store [56]. As LGSVL is a Unity-based simulator, these could be directly placed into scenarios. The roadside objects we utilised in TrashFuzz have been assembled into a single library that we have made available online [51]. To ensure the objects are placed according to conventional road design guidelines, we constrained their placement according to the functions in Table I. These functions take a scenario as input and return a Boolean value to indicate whether the generated scenario is valid or not according to the guidelines.
Finally, we integrated the above components and implemented our greedy gradient-based fuzzing engine as described in Section III-B. Our experiments were conducted using two machines, each equipped with 32GB of memory, an Intel i7-10700k CPU, and an RTX 2080Ti graphics card. These two machines run on Linux (Ubuntu 20.04.5 LTS) and Windows (Windows 10 Pro) operating systems, respectively.
IV-B Evaluation
We conducted experiments to answer four Research Questions (RQs):
  • RQ1: Can TrashFuzz produce valid scenarios?